OpenStudy (anonymous):

Let A = total area of five circles of radius r and let B = total area of three circles of radius s. If A = B, then r / s =

OpenStudy (anonymous):

\[ A = 5 \pi r^2 \\ B = 2 \pi s^2 \\ A = B \implies 5 \pi r^2 = 2 \pi s^2 \]

OpenStudy (anonymous):

Can you solve for \(r/s\)?

OpenStudy (watchmath):

@wio: \(B=3\pi s^2\)